OpenStudy (anonymous):

Can anyone explain to me how to write the equation of the line that passes through (2, 4) and (1, –3) in slope-intercept form ?

OpenStudy (anonymous):

First find the slope (-3-4)/(1-2). Then pick one of the points (2,4) might be easier because they're both positive. Then plug your numbers into the point-slope formula, y-y1=m(x-x1). Do you know how to do that?

OpenStudy (anonymous):

I think so,, so i would have to solve 4(-3) = 7 (2-1) ?

OpenStudy (anonymous):

where's the y's and x's in your equation?

OpenStudy (anonymous):

y-(-3) = 7 (x - 1) ?

OpenStudy (anonymous):

yep... just simplify it...

OpenStudy (anonymous):

good job - you got it from there?

OpenStudy (anonymous):

7=7?... No, I don't lol and i feel like an idiot - I dont see how I'm supposed to get it in slope intercept form ..

OpenStudy (anonymous):

when you simplify what you got above, do you get y+3=7x-7?

OpenStudy (anonymous):

If you do, subtract 3 from both sides and you'll have it in slope-intercept form!

OpenStudy (anonymous):

ohhhh y = 7x – 10 ? Thank you!!

OpenStudy (anonymous):

that's it - works like a charm! (You're welcome)